OpenStudy (anonymous):

In triangle ABC, what is the measure of A if a = 15, b = 10, and c = 7? 57.12° 72.15° 103.21° 122.88°

OpenStudy (anonymous):

for this one, use the law of cosines so, a^2 = b^2 + c^2 - 2(b)(c)cosA. for you, thats: 15^2 = 10^2 + 7^2 - 2(10)(7)cosA

OpenStudy (anonymous):

225 = 149 - 140cosA. subtract the 149 from both sides so you have 76 = -140cosA

OpenStudy (anonymous):

divide 76 by -140

OpenStudy (anonymous):

the answer is not an answer choice

OpenStudy (anonymous):

you should get a pretty long decimal that is about -0.5429

OpenStudy (anonymous):

take the inverse cosine of that decimal

OpenStudy (anonymous):

your answer is about 122.88 and that would be degrees

OpenStudy (anonymous):

ohh i understand now!! thank you!!

OpenStudy (anonymous):

sure thing